OpenStudy (anonymous):

Calculate the boiling point elevation of a solution containing 69.1 g of glucose (C6H12O6) dissolved in 380.6 g of water. Calculate the freezing point depression for the same solution. boiling point elevation

OpenStudy (aaronq):

use the formula: \(\Delta T=i*m*K_{\theta}\) where, \(\theta=b\) when finding BP, and \(\theta=f\) when finding freezing point (By that, i mean, use the appropriate constant.) More on this: http://openstudy.com/study#/updates/521ecb3ae4b0750826e0c362